OpenStudy (anonymous):

A ball rolling downhill hits a top speed of 30 mph. Which speeds are the same as the ball’s top speed? Choose all answers that are correct. (Points : 3) 720 mi/day 1800 mi/min 158,400 ft/h 9,504,000 ft/min

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

30 miles per hour = 30*24 miles per day = 720 mi/day

OpenStudy (anonymous):

only a is correct then

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

anyone there

OpenStudy (aum):

1 mile = 5280 feet 30 miles / hour = 30 * 5280 feet / hour = 158,400 ft / h

OpenStudy (math&ing001):

A and C @Xgloves11

OpenStudy (anonymous):

thanks guys